OpenStudy (kayleigh98):

Write the equation of the line that goes through the point (−3, −5) and is parallel to the line −x + 3y = −0.5.

OpenStudy (lynfran):

when 2 lines are parallel the have the same slope so -x+3y=-0.5 3y=x-0.5 y=1/3(x)- (0.5)/3 slope =1/3 now we use the given point (-3,-5) with the slope to find such equation...

OpenStudy (lynfran):

here the formula y=mx+b where y=-5 , x=-3 and m=1/3 subs those values to find b...its y intercept then....

OpenStudy (kayleigh98):

b=-6?

OpenStudy (lynfran):

-5=1/3(-3)+b -5=-1+b -5+1=b 1-5=b

OpenStudy (kayleigh98):

-4?

OpenStudy (lynfran):

yep

OpenStudy (lynfran):

so y=mx+b we now sub our "m" and "b" into the formula

OpenStudy (lynfran):

so our parallel equation is??

OpenStudy (kayleigh98):

y=1/3x-4

OpenStudy (lynfran):

y=1/3(x)-4 can also be written as 3y=x-12 3y-x=-12 yep